# Stable Differentiable Modal Synthesis for Learning Nonlinear Dynamics

### v1

@@ Abstract
Modal methods are a long-standing approach to physical modelling synthesis. Extensions to nonlinear problems are possible, including the case of a high-amplitude vibration of a string. A modal decomposition leads to a densely coupled nonlinear system of ordinary differential equations. Recent work in scalar auxiliary variable techniques has enabled construction of explicit and stable numerical solvers for such classes of nonlinear systems. On the other hand, machine learning approaches (in particular neural ordinary differential equations) have been successful in modelling nonlinear systems automatically from data. In this work, we examine how scalar auxiliary variable techniques can be combined with neural ordinary differential equations to yield a stable differentiable model capable of learning nonlinear dynamics. The proposed approach leverages the analytical solution for linear vibration of system’s modes so that physical parameters of a system remain easily accessible after the training without the need for a parameter encoder in the model architecture. As a proof of concept, we generate synthetic data for the nonlinear transverse vibration of a string and show that the model can be trained to reproduce the nonlinear dynamics of the system. Sound examples are presented.
@@ 1  INTRODUCTION
In this work, we aim to more deeply integrate the physics of a distributed system into a machine learning framework. We use modal decomposition to construct a system of finite dimension and separate the linear and nonlinear parts of the problem. Then, we replace only a dimensionless memoryless nonlinearity that describes coupling between the modes with a neural network. In particular, we employ gradient networks (GradNets) [5] that allow an interpretation in terms of a closed-form and non-negative potential function. Finally, we use neural ordinary differential equations (NODEs) [6] in combination with the SAV technique to yield a stable differentiable model that can learn nonlinear dynamics from data. As a result, physical parameters remain easily accessible and the model generalises to physical parameters, sampling rates and time scales not seen during training.
This paper is an expanded version of a paper presented at DAFx25 [36] and is organised as follows. A simple model of nonlinear transverse string vibration is described in Section 2. Section 3 derives modal equations for a string, which are then discretised in time using an explicit and stable numerical solver in Section 4. Section 5 outlines the proposed differentiable model, which is evaluated in Section 6. Sound examples are available on the accompanying page [^1] https://victorzheleznov.github.io/jaes-modal-node.
@@ 2  NONLINEAR TRANSVERSE STRING VIBRATION
$u(0,t)=\partial_{x}^{2}u(0,t)=u(L,t)=\partial_{x}^{2}u(L,t)=0,\;\forall t\in\mathbb{R}^{+}.$
Output is assumed to be drawn directly from the string displacement at position $x_{\mathrm{o}}$ as $w(t)=u(x_{\mathrm{o}},t)$.
@@ 2.1  Linear Vibration
where $\partial_{t}$ and $\partial_{x}$ represent partial derivatives with respect to $x$ and $t$, respectively. Physical parameters that appear in $\mathcal{L}$ are: the material density $\rho$ in $\mathrm{kg{\cdot}m^{-3}}$; the string cross-sectional area $A=\pi r^{2}$ in $\mathrm{m^{2}}$ for a string of radius $r$; the tension $T_{0}$ in $\mathrm{N}$; Young’s modulus $E$ in $\mathrm{N{\cdot}m^{-2}}$; and moment of inertia $I=\frac{1}{4}\pi r^{4}$ in $\mathrm{m^{4}}$. Frequency-independent and dependent loss is characterised by parameters $\sigma_{0}\geq 0$ and $\sigma_{1}\geq 0$, respectively. See [3] for more on these terms in the context of linear string vibration.
@@ 2.2  Nonlinearity
$\mathcal{F}(x,t)=\frac{EA-T_{0}}{2}\partial_{x}\big(\mathcal{V}^{\prime}(\xi)\big),\quad\xi\triangleq\partial_{x}u,$
@@ 2.4  Equation Scaling
where $\gamma=\frac{1}{L}\sqrt{\frac{T_{0}}{\rho A}}$, $\alpha=\sqrt{\frac{EA}{T_{0}}}$, $\kappa=\frac{1}{L^{2}}\sqrt{\frac{EI}{\rho A}}$, $\hat{\sigma}_{1}=\frac{\sigma_{1}}{L^{2}}$ and $\hat{f}_{\mathrm{e}}(t)=\frac{1}{\rho AL^{2}}f_{\mathrm{e}}(t)$.
Thus, we have reduced a set of physical parameters $\{L,\rho,A,\allowbreak T_{0},E,I,\sigma_{0},\sigma_{1}\}$ to a set of only five parameters $\{\gamma,\kappa,\alpha,\sigma_{0},\hat{\sigma}_{1}\}$. In the following sections we omit the “hat” while referring to the scaled string model (4).
@@ 3  MODAL DECOMPOSITION
The solution to equation (4) can be decomposed into a set of modes, yielding a finite-dimensional system when truncated to finite order $M$. The transverse displacement $u$ is rewritten as a superposition of modal displacements $\mathbf{q}(t)=[q_{1}(t),\dots,q_{M}(t)]^{\mathrm{T}}$:
@@
Substituting (3) into (4), left-multiplying by $\bm{\Phi}$ and taking an $L^{2}$ inner product over the interval $[0,1]$, we obtain the following second-order system of ODEs:
@@
The system (3) can be rewritten in the first-order form using modal velocities $\mathbf{p}(t)=[p_{1}(t),\dots,p_{M}(t)]^{\mathrm{T}}$ as:
@@ 3.1  Spectral Method
To obtain a computationally efficient expression for the dimensionless nonlinearity $\mathbf{f}(\mathbf{q})\!\!:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$, we employ a spectral method [35, 34] to calculate spatial derivatives $\partial_{x}$ on the grid $x_{l+\nicefrac{{1}}{{2}}}=\frac{1}{M+1}\big(l+\frac{1}{2}\big),\;l=0,\dots,M$.
@@ 4.1  Quadratisation
To guarantee stability of numerical simulation, we make use of the SAV technique [2, 32]. SAV methods are based on quadratisation of the potential $V(\mathbf{q})$ through transformation $\psi\triangleq\sqrt{2V(\mathbf{q})+\epsilon}$ where $\epsilon>0$ is a gauge constant. This transformation is well defined if the potential $V(\mathbf{q})$ (3.1) is non-negative which is the case for the chosen nonlinear function (2.2).
@@ 5.1  Neural Ordinary Differential Equations
Here $\mathbf{y}=\mathbf{y}(t)\!\!:\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is an unknown function of time $t$, $\mathbf{y}_{0}\in\mathbb{R}^{K}$ is an initial condition and $\mathbf{h}_{\theta}(\mathbf{y},t)\!\!:\mathbb{R}^{K}\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is a neural network where $\theta$ denotes the set of all learnable parameters and $K$ denotes the state dimension. Generally, a simple architecture such as a multilayer perceptron is chosen for $\mathbf{h}_{\theta}(\mathbf{y},t)$. Chen et al. [6] have showed that the system (5.1) in combination with a numerical solver, labelled as ODENet, can be trained from data to reproduce dynamics of a target system.
@@ 5.2  Extension for Modal Synthesis
The formulation (5.2) has strong implications. The exact expression for linear vibration exploits the periodic, harmonic and lossy nature of a musical system. Thus, we aid the optimisation process by constraining the space of possible solutions and improve interpretability of the model. Furthermore, the neural network $\mathbf{f}_{\theta}(\mathbf{q})$ is memoryless and dimensionless, thus does not depend on physical parameters of a system and external excitation. Theoretically, these parameters can be changed after the training as long as range of displacements $\mathbf{q}$ stays the same as in a training dataset to simulate other configurations of a system.
@@ 5.3  Gradient Networks
The nature of the SAV technique (Section 4.1) constrains possible network architectures for $\mathbf{f}_{\theta}(\mathbf{q})$ as we require existence of a closed-form and non-negative potential $V_{\theta}(\mathbf{q})$. To overcome this, we employ GradNets [5] that directly parametrise gradients of various function classes. In particular, we use the following architecture:
@@ 6  EVALUATION
For the parametrisation of $\mathbf{f}_{\theta}(\mathbf{q})$ we used the GradNet (5.3) with hidden dimension $H=1000$. A leaky rectified linear unit (Leaky ReLU) was used as an activation $\sigma$ which is a monotonically-increasing function with a non-negative antiderivative $\phi$. Kaiming initialisation [12] was used for the initial weights $\mathbf{W}$ and the biases $\mathbf{b}$ were initialised to zero. Logarithms of scaling vectors $\bm{\alpha}$ and $\bm{\beta}$ were used as learnable parameters to enforce non-negativity and were initialised around zero using a normal distribution.
The training loss was the MSE (5.1) taken over the state vector $\mathbf{y}^{\mathrm{T}}=[\mathbf{q}^{\mathrm{T}},\mathbf{p}^{\mathrm{T}}]$, excluding the auxiliary variable $\psi$ arising from the SAV method. This implied that the potential $V_{\theta}(\mathbf{q})$ could be shifted by a constant relative to the target potential (3.1) since dynamics of the system and the training loss would not be affected. Backpropagation was performed using internal operations of the numerical solver (4.2), i.e., the “discretise-then-optimise” method, which is generally a preferred approach due to its gradient accuracy, speed and straightforward implementation [13]. Calculation of the control term $\mathbf{g}_{\mathrm{mod}}$ was excluded from the computational graph as it was found to significantly compromise training by steering the potential $V_{\theta}(\mathbf{q})$ to incorrect values of the auxiliary variable $\psi$. The Adam optimiser [14] was used with default parameters. Three separate datasets for training, validation and testing were used (Section 6.1). The training was performed for 2000 epochs using a cluster of four GPUs to parallelise batches of data. The resulting model was chosen based upon the lowest loss obtained on the validation dataset.
@@ 6.1  Datasets
Three datasets were independently generated using the nonlinear transverse string model (Section 2). The training and test datasets both consisted of 60 trajectories and the validation dataset consisted of 20 trajectories. Each trajectory included both displacement and velocity information for each mode of a string. The number of modes was equal to $M=75$. Taking into account the effect of stiffness, this covered the $\mathrm{10}$$\mathrm{kHz}$ range for the lowest considered fundamental frequency at $\mathrm{61{.}74}$$\mathrm{Hz}$. Oversampling by a factor of two for standard sampling rates $\mathrm{44{.}1}$$\mathrm{kHz}$ and $\mathrm{48}$$\mathrm{kHz}$ was used to avoid aliasing due to the nonlinear effects for high amplitude excitations.
Simulation parameters used for the datasets are provided in Table 1, where $T_{\mathrm{sim}}$ corresponds to the duration of simulation. For specified parameter ranges, randomised values were generated from a uniform distribution. For training, fundamental frequencies, equal to $\frac{\gamma}{2}$, spanned a half-octave range from $\mathrm{61{.}74}$$\mathrm{Hz}$ (B1 note) to $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note). For validation and testing, the next half-octave range from $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note) to $\mathrm{123{.}47}$$\mathrm{Hz}$ (B2 note) was used. In addition, the simulation duration was increased to account for a longer decay time due to a smaller damping parameter $\sigma_{0}$. Non-overlapping ranges for the stiffness parameter $\kappa$ were used. Strings were excited by randomly-generated excitation functions (2.3) at randomised excitation positions $x_{\mathrm{e}}$. The range for the excitation amplitude $f_{\mathrm{amp}}$ was scaled linearly with fundamental frequencies to preserve the strength of the nonlinear effects between the datasets, and thus the range of displacements $\mathbf{q}$ (see Appendix A.1). An audio output was drawn from randomised positions $x_{\mathrm{o}}$ along a string for each trajectory.
@@ 6.2  Results
For evaluation of the trained model, we used the relative MSE, defined for a general time series $\mathbf{x}^{n}$ as:
@@
Metrics for predicted displacement trajectories $\tilde{\mathbf{q}}^{n}$ and audio outputs $\tilde{w}^{n}$, averaged over the training, validation and test datasets, are provided in Table 2. Metrics were separately computed for the initial $\mathrm{100}$$\mathrm{ms}$ and for the full duration of simulation. As can be seen, metrics for the validation and test datasets have the same order of magnitude as for the training dataset and do not degrade for unseen simulation parameters. This suggests that the sampling rate and duration of simulation can be changed after the training and the model generalises to unseen physical parameters as long as the range of displacements $\mathbf{q}^{n}$ stays the same.
@@
Figure 1: The predicted displacement trajectory with the largest $\operatorname{MSE}_{\mathrm{rel}}(\tilde{w}^{n},w^{n})$ (centre), the target displacement trajectory (left) and the relative absolute error between them (right).
@@
Formal perceptual tests are out of scope of this paper and readers are encouraged to listen to sound examples presented on the accompanying page [^3] https://victorzheleznov.github.io/jaes-modal-node.

### v2

@@ Abstract
Modal methods are a long-standing approach to physical modelling synthesis. Extensions to nonlinear problems are possible, leading to coupled nonlinear systems of ordinary differential equations. Recent work in scalar auxiliary variable techniques has enabled construction of explicit and stable numerical solvers for such systems. On the other hand, neural ordinary differential equations have been successful in modelling nonlinear systems from data. In this work, we examine how scalar auxiliary variable techniques can be combined with neural ordinary differential equations to yield a stable differentiable model capable of learning nonlinear dynamics. The proposed approach leverages the analytical solution for linear vibration of the system’s modes so that physical parameters of a system remain easily accessible after the training without the need for a parameter encoder in the model architecture. Compared to our previous work that used multilayer perceptrons to parametrise nonlinear dynamics, we employ gradient networks that allow an interpretation in terms of a closed-form and non-negative potential required by scalar auxiliary variable techniques. As a proof of concept, we generate synthetic data for the nonlinear transverse vibration of a string and show that the model can be trained to reproduce the nonlinear dynamics of the system. Sound examples are presented.
@@ 1  INTRODUCTION
In this paper we expand our physics-informed machine learning framework presented at DAFx25 conference [38] by introducing stability guarantees. We use modal decomposition to construct a system of finite dimension, separate the linear and nonlinear parts of the problem and replace only a dimensionless memoryless nonlinearity that describes coupling between the modes with a neural network. In contrast to our previous work [38] that relied on multilayer perceptrons (MLPs), we impose additional architectural constraints [23] through gradient networks (GradNets) [5] that allow an interpretation in terms of a closed-form and non-negative potential function. This enables application of the SAV technique for numerical integration of neural ordinary differential equations (NODEs) [6], leading to a stable differentiable model that can learn nonlinear dynamics from data. As a result of separating the linear vibration of a system from the neural network, physical parameters remain easily accessible and the model generalises to physical parameters, sampling rates and time scales not seen during training.
The paper is organised as follows. A simple model of nonlinear transverse string vibration is described in Section 2. Section 3 derives modal equations for a string, which are then discretised in time using an explicit and stable numerical solver in Section 4. Section 5 outlines the proposed differentiable model based on the obtained modal equations, which is evaluated for the case study of nonlinear transverse string vibration in Section 6. Sound examples are available on the accompanying page [^1] https://victorzheleznov.github.io/jaes-modal-node.
@@ 2  NONLINEAR TRANSVERSE STRING VIBRATION
$u(0,t)=\partial_{x}^{2}u(0,t)=u(L,t)=\partial_{x}^{2}u(L,t)=0,\;\forall t\in\mathbb{R}^{+},$
where $\partial_{x}$ represents a partial derivative with respect to $x$. Output is assumed to be drawn directly from the string displacement at position $x_{\mathrm{o}}$ as $w(t)=u(x_{\mathrm{o}},t)$.
@@ 2.1  Linear Vibration
where $\partial_{t}$ represents a partial derivative with respect to $t$. Physical parameters that appear in $\mathcal{L}$ are: the material density $\rho$ in $\mathrm{kg{\cdot}m^{-3}}$; the string cross-sectional area $A=\pi r^{2}$ in $\mathrm{m^{2}}$ for a string of radius $r$; the tension $T_{0}$ in $\mathrm{N}$; Young’s modulus $E$ in $\mathrm{N{\cdot}m^{-2}}$; and moment of inertia $I=\frac{1}{4}\pi r^{4}$ in $\mathrm{m^{4}}$. Frequency-independent and dependent loss is characterised by parameters $\sigma_{0}\geq 0$ and $\sigma_{1}\geq 0$, respectively. See [3] for more on these terms in the context of linear string vibration.
@@ 2.2  Nonlinearity
$\mathcal{F}(x,t)=\frac{EA-T_{0}}{2}\partial_{x}\big(\mathcal{V}^{\prime}(\xi)\big),\quad\xi=\xi(x,t)\triangleq\partial_{x}u,$
@@ 2.4  Equation Scaling
where $\gamma=\frac{1}{L}\sqrt{\frac{T_{0}}{\rho A}}$, $\kappa=\frac{1}{L^{2}}\sqrt{\frac{EI}{\rho A}}$, $\alpha=\sqrt{\frac{EA}{T_{0}}}$, $\hat{\sigma}_{1}=\frac{\sigma_{1}}{L^{2}}$ and $\hat{f}_{\mathrm{e}}(t)=\frac{1}{\rho AL^{2}}f_{\mathrm{e}}(t)$.
Thus, we have reduced a set of physical parameters $\{L,\rho,A,\allowbreak T_{0},E,I,\sigma_{0},\sigma_{1}\}$ to a set of only five parameters $\{\gamma,\kappa,\alpha,\sigma_{0},\hat{\sigma}_{1}\}$. In the following sections we omit the “hat” while referring to the scaled string model (4), including the loss parameter $\hat{\sigma}_{1}$ and the excitation function $\hat{f}_{\mathrm{e}}(t)$.
@@ 3  MODAL DECOMPOSITION
The solution to equation (4) can be decomposed into a set of modes [3], yielding a finite-dimensional system when truncated to finite order $M$. The transverse displacement $u$ is rewritten as a superposition of modal displacements $\mathbf{q}=\mathbf{q}(t)=[q_{1}(t),\dots,q_{M}(t)]^{\mathrm{T}}$:
@@
Substituting (3) into (4), left-multiplying by $\bm{\Phi}(x)$ and taking an $L^{2}$ inner product over the interval $[0,1]$, we obtain the following second-order system of ODEs:
@@
The system (3) can be rewritten in the first-order form using modal velocities $\mathbf{p}=\mathbf{p}(t)=[p_{1}(t),\dots,p_{M}(t)]^{\mathrm{T}}$ as:
@@ 3.1  Spectral Method
To obtain a closed-form expression for the dimensionless nonlinearity $\mathbf{f}(\mathbf{q})\!\!:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$, we employ a spectral method [37, 36] to calculate spatial derivatives $\partial_{x}$ on the grid $x_{l+\nicefrac{{1}}{{2}}}=\frac{1}{M+1}\big(l+\frac{1}{2}\big),\;l=0,\dots,M$. Compared to our previous work [38] that calculated spatial derivatives $\partial_{x}$ over a continuous domain, the spectral method results in a significantly more efficient expression for $\mathbf{f}(\mathbf{q})$ and allows to implement the nonlinear function (2.2) without approximation by Taylor series.
@@ 4.1  Quadratisation
In our previous work [38], we employed the Störmer-Verlet numerical solver which does not guarantee stability in simulation. To address this, we make use of the SAV technique [2, 34]. We introduce an auxiliary variable $\psi$ in order to quadratise the potential $V(\mathbf{q})$:
$\psi\triangleq\sqrt{2V(\mathbf{q})+\epsilon},$
where $\epsilon>0$ is a gauge constant. The auxiliary variable $\psi$ is well defined if the potential $V(\mathbf{q})$ (3.1) is non-negative which is the case for the chosen nonlinear function (2.2).
@@ 5.1  Neural Ordinary Differential Equations
Here $\mathbf{y}=\mathbf{y}(t)\!\!:\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is an unknown function of time $t$, $\mathbf{y}_{0}\in\mathbb{R}^{K}$ is an initial condition and $\mathbf{h}_{\theta}(\mathbf{y},t)\!\!:\mathbb{R}^{K}\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is a neural network where $\theta$ denotes the set of all learnable parameters and $K$ denotes the state dimension. Generally, a simple architecture such as an MLP is chosen for $\mathbf{h}_{\theta}(\mathbf{y},t)$. Chen et al. [6] have showed that the system (5.1) in combination with a numerical solver, labelled as ODENet, can be trained from the observed state data to reproduce the dynamics of a target system for which the theoretical model may be unknown.
@@ 5.2  Extension for Modal Synthesis
The formulation (5.2) has strong implications. The exact expression for linear vibration exploits the periodic, harmonic and lossy nature of a musical system. Thus, we aid the optimisation process by constraining the space of possible solutions and improve interpretability of the model. Furthermore, the neural network $\mathbf{f}_{\theta}(\mathbf{q})$ is memoryless and dimensionless, thus does not depend on physical parameters of a system and external excitation. Theoretically, these parameters can be changed after the training as long as range of displacements $\mathbf{q}$ stays the same as in a training dataset to simulate other configurations of a system. Some limitations exist since the neural network $\mathbf{f}_{\theta}(\mathbf{q})$ implicitly depends on boundary conditions and the number of modes.
@@ 5.3  Gradient Networks
The nature of the SAV technique (Section 4.1) constrains possible network architectures for $\mathbf{f}_{\theta}(\mathbf{q})$ as we require existence of a closed-form and non-negative potential $V_{\theta}(\mathbf{q})$. For example, we can no longer use MLPs as in our previous work [38]. To overcome this, we employ GradNets [5] that directly parametrise gradients of various function classes. In particular, we use the following architecture:
@@ 6  EVALUATION
For the parametrisation of $\mathbf{f}_{\theta}(\mathbf{q})$ we used the GradNet (5.3) with hidden dimension $H=1000$. This number was chosen based on available GPU memory since the validation loss was found to consistently decrease with increasing hidden dimension. However, smaller networks may also be sufficient from the perceptual point of view. A leaky rectified linear unit was used as an activation $\sigma$ which is a monotonically-increasing function with a non-negative antiderivative $\phi$. Kaiming initialisation [12] was used for the initial weights $\mathbf{W}$ and the biases $\mathbf{b}$ were initialised to zero. Logarithms of scaling vectors $\bm{\alpha}$ and $\bm{\beta}$ were used as learnable parameters to enforce non-negativity and were initialised around zero using a normal distribution.
The training loss was the MSE (5.1) taken over the state vector $\mathbf{y}^{\mathrm{T}}=[\mathbf{q}^{\mathrm{T}},\mathbf{p}^{\mathrm{T}}]$, excluding the auxiliary variable $\psi$ arising from the SAV method. This implied that the potential $V_{\theta}(\mathbf{q})$ could be shifted by a constant relative to the target potential (3.1) since dynamics of the system and the training loss would not be affected. Backpropagation was performed using internal operations of the numerical solver (4.2), i.e., the “discretise-then-optimise” method, which is generally a preferred approach due to its gradient accuracy, speed and straightforward implementation [13]. Calculation of the control term $\mathbf{g}_{\mathrm{mod}}$ was excluded from the computational graph. In some cases, it was found to significantly compromise training by steering the potential $V_{\theta}(\mathbf{q})$ to incorrect values of the auxiliary variable $\psi$ and lead to an unexpected increase in the training loss. As only gradient calculation is affected, stability and drift regulation properties of the numerical solver (4.2) still hold during inference. The Adam optimiser [14] was used with default parameters. The training was performed for 2000 epochs using a cluster of four GPUs to parallelise batches of data. The resulting model was chosen based on the lowest validation loss obtained during optimisation.
@@ 6.1  Datasets
Three separate datasets for training, validation and testing were independently generated using the nonlinear transverse string model (Section 2). The training and test datasets both consisted of 60 trajectories and the validation dataset consisted of 20 trajectories. Each trajectory included both displacement and velocity information for each mode of a string. The number of modes was equal to $M=75$. Taking into account the effect of stiffness, this covered the $\mathrm{10}$$\mathrm{kHz}$ range for the lowest considered fundamental frequency at $\mathrm{61{.}74}$$\mathrm{Hz}$. Oversampling by a factor of two for standard sampling rates $\mathrm{44{.}1}$$\mathrm{kHz}$ and $\mathrm{48}$$\mathrm{kHz}$ was used to avoid aliasing due to the nonlinear effects for high amplitude excitations.
@@
Simulation parameters used for the datasets are provided in Table 1, where $T_{\mathrm{sim}}$ corresponds to the duration of simulation. For specified parameter ranges, randomised values were generated from a uniform distribution. For training, fundamental frequencies, equal to $\frac{\gamma}{2}$, spanned a half-octave range from $\mathrm{61{.}74}$$\mathrm{Hz}$ (B1 note) to $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note). For validation and testing, the next half-octave range from $\mathrm{87{.}31}$$\mathrm{Hz}$ (F2 note) to $\mathrm{123{.}47}$$\mathrm{Hz}$ (B2 note) was used. In addition, the simulation duration was increased to account for a longer decay time due to a smaller loss parameter $\sigma_{0}$. Non-overlapping ranges for the stiffness parameter $\kappa$ were used. Strings were excited by randomly-generated excitation functions (2.3) at randomised excitation positions $x_{\mathrm{e}}$. The range for the excitation amplitude $f_{\mathrm{amp}}$ was scaled linearly with fundamental frequencies to preserve the strength of the nonlinear effects between the datasets, and thus the range of displacements $\mathbf{q}$ (see Appendix A.1). An audio output was drawn from randomised positions $x_{\mathrm{o}}$ along a string for each trajectory.
@@ 6.2  Results
Figure 1: The predicted displacement trajectory with the largest $\operatorname{MSE}_{\mathrm{rel}}(\tilde{w}^{n},w^{n})$ (centre), the target displacement trajectory (left) and the pointwise absolute error between them relative to the maximum absolute value of target trajectory (right).
For evaluation of the trained model, we used the relative MSE and the relative mean absolute error (MAE), defined for a general time series $\mathbf{x}^{n}$ as:
@@
$\operatorname{MAE}_{\mathrm{rel}}(\tilde{\mathbf{x}}^{n},\mathbf{x}^{n})\triangleq\frac{\sum_{n}\left\lVert\tilde{\mathbf{x}}^{n}-\mathbf{x}^{n}\right\rVert_{1}}{\sum_{n}\left\lVert\mathbf{x}^{n}\right\rVert_{1}},$
@@
Metrics for predicted displacement trajectories $\tilde{\mathbf{q}}^{n}$ and audio outputs $\tilde{w}^{n}$, averaged over the training, validation and test datasets, are provided in Table 2. Metrics were separately computed for the initial $\mathrm{100}$$\mathrm{ms}$ and for the full duration of simulation. As can be seen, metrics for the validation and test datasets have the same order of magnitude as for the training dataset and do not degrade for unseen simulation parameters. This suggests that the sampling rate and duration of simulation can be changed after the training and the model generalises to unseen physical parameters as long as the range of displacements $\mathbf{q}^{n}$ stays the same. In addition, values of the relative MAE are close to the square roots of the relative MSE, indicating the absence of outliers in the datasets.
@@
$\operatorname{MAE}_{\mathrm{rel}}(\tilde{\mathbf{q}}^{n},\mathbf{q}^{n})$ $\mathrm{4{.}1{}\times{}10^{-2}}$ $\mathrm{3{.}4{}\times{}10^{-2}}$ $\mathrm{3{.}4{}\times{}10^{-2}}$
$\operatorname{MAE}_{\mathrm{rel}}(\tilde{w}^{n},w^{n})$ $\mathrm{1{.}6{}\times{}10^{-2}}$ $\mathrm{1{.}1{}\times{}10^{-2}}$ $\mathrm{1{.}3{}\times{}10^{-2}}$
@@
$\operatorname{MAE}_{\mathrm{rel}}(\tilde{\mathbf{q}}^{n},\mathbf{q}^{n})$ $\mathrm{3{.}6{}\times{}10^{-1}}$ $\mathrm{3{.}9{}\times{}10^{-1}}$ $\mathrm{3{.}9{}\times{}10^{-1}}$
$\operatorname{MAE}_{\mathrm{rel}}(\tilde{w}^{n},w^{n})$ $\mathrm{3{.}1{}\times{}10^{-1}}$ $\mathrm{3{.}3{}\times{}10^{-1}}$ $\mathrm{3{.}5{}\times{}10^{-1}}$
@@
Formal perceptual tests are out of scope of this paper and readers are encouraged to listen to sound examples presented on the accompanying page [^3] https://victorzheleznov.github.io/jaes-modal-node. Based on informal listening, the predicted and target audio were found to be nearly indistinguishable, while the difference compared to the linear baseline was clearly audible.